# Refereed Learning

### v1

@@ 7 Protocols with additive and mixed error
First, we briefly consider additive error in the single prover setting. That is, given query access to $h_{0}$, $h_{1}$, and $f$, a verifier $\mathcal{V}$, with the help of a prover $\mathcal{P}$}, would like to decide which of $h_{0}$ and $h_{1}$ has better loss on $f$. Prior work of [GRSY21] gives a protocol for empirical risk minimization in the single prover setting that can be easily adapted to compare $h_{0}$ and $h_{1}$. At a high level, the protocol works as follows: [^13] We describe a modified version of their protocol tailored to our setting. For a complete description of the protocol see the proof of Claim 5.2 (Simple Query Delegation) in [GRSY21]. (1) The verifier draws $\Theta{\left({\frac{1}{\eta^{2}}}\right)}$ unlabeled samples from $\mathcal{D}$} and $\Theta{\left({\frac{1}{\eta}}\right)}$ labeled samples $(x,f(x))$ where $x\sim{\mathcal{D}}$, and sends the samples (not including labels) to the prover. (2) The prover labels the samples using $f$ and sends them back to the verifier. (3) The verifier checks that the prover’s labels agree with its own set of labeled samples. If the labels disagree then the verifier rejects. Otherwise, the verifier uses the labels, along with query access to $h_{0}$ and $h_{1}$, to determine which of $h_{0}$ and $h_{1}$ achieves smaller loss on $f$.
@@ Acknowledgments
We thank Mark Bun for simplifying the proof of 4.6.

### v2

@@ 7 Protocols with additive and mixed error
First, we briefly consider additive error in the single prover setting. That is, given query access to $h_{0}$, $h_{1}$, and $f$, a verifier $\mathcal{V}$, with the help of a prover $\mathcal{P}$}, would like to decide which of $h_{0}$ and $h_{1}$ has better loss on $f$. Prior work of [GoldwasserRSY21] gives a protocol for empirical risk minimization in the single prover setting that can be easily adapted to compare $h_{0}$ and $h_{1}$. At a high level, the protocol works as follows: [^14] We describe a modified version of their protocol tailored to our setting. For a complete description of the protocol see the proof of Claim 5.2 (Simple Query Delegation) in [GoldwasserRSY21]. (1) The verifier draws $\Theta{\left({\frac{1}{\eta^{2}}}\right)}$ unlabeled samples from $\mathcal{D}$} and $\Theta{\left({\frac{1}{\eta}}\right)}$ labeled samples $(x,f(x))$ where $x\sim{\mathcal{D}}$, and sends the samples (not including labels) to the prover. (2) The prover labels the samples using $f$ and sends them back to the verifier. (3) The verifier checks that the prover’s labels agree with its own set of labeled samples. If the labels disagree then the verifier rejects. Otherwise, the verifier uses the labels, along with query access to $h_{0}$ and $h_{1}$, to determine which of $h_{0}$ and $h_{1}$ achieves smaller loss on $f$.
@@ Acknowledgments
We thank Mark Bun for simplifying the proof of 4.6. We thank Noam Mazor for significantly simplifying the certifiable sample protocol (Lemma 3.1).